### v1

# Emergence of non-ergodic multifractal quantum states in geometrical fractals

Abstract
Eigenstate multifractality, a hallmark of non-interacting disordered metals which can potentially be observed in many-body localized states as well, is characterized by anomalous slow dynamics and appears relevant for many areas of quantum physics from measurement-driven systems to superconductivity. We propose a novel approach to achieve non-ergodic multifractal (NEM) states without disorder by iteratively introducing defects into a crystal lattice, reshaping it from a plain structure into fractal geometry. By comprehensive analysis of the Sierpiński gasket case, we find a robust evidence of the emergence of NEM states that go beyond the conventional classification of quantum states and designate new pathways for quantum transport studies. We discuss potential experimental signatures of these states.
Introduction
Since the discovery of Anderson’s localization paradigm [1], there has been a significant interest in critical phenomena that transcend the traditional dichotomy of extended and localized states. Among these, a class of novel states, commonly referred to as non-ergodic multifractal states (NEMs) [2, 3, 4], has attracted particular attention. NEMs occupy a finite volume of the system but their support is a set of points that, in the thermodynamic limit, has a zero measure, challenging the conventional ideas of how a quantum state can be distributed in space. The concept of multifractality finds broad applications in a wide range of domains of science and engineering, including fluid dynamics [5], topography [6], seismology [7], finance [8], physiology [9], and epidemic modeling [10]. It also inspired advancements in machine learning, granting unexplored settings to incorporate larger hierarchical depth and achieve stable and robust prediction performances [11, 12]. In condensed matter physics, the concept of NEM is strictly connected to the singular continuous nature of the spectrum [13, 14, 15] and the critical scaling behavior [16, 17, 18] observed at phase transitions, which give rise to anomalous subdiffusive dynamics [19, 20, 21] and make NEM particularly intriguing in the context of quantum transport. Additionally, they exhibit sublinear growth of entanglement entropy [22, 23], defying the standard formulation of eigenstate thermalization hypothesis [24]. These attributes are essential for understanding quantum coherence of eigenstates [25] and many-body localization phenomena [26, 27]. Moreover, it is conjectured that multifractality can enhance electron-electron interactions leading to elevated critical temperature in superconductors [28, 29].
Incommensurate potentials in one-dimensional systems [30, 31, 32, 33] and randomness in higher-dimensional systems [34, 35, 36, 37] are genuine mechanisms to induce disorder causing the emergence of NEMs in quantum systems. However, their direct experimental observations remain problematic due to finite-size effects and the dependence on fine-tuning, posing considerable difficulties for further exploration of NEMs. In this work, we propose an approach to generating such states. Rather than relying on randomness to induce disorder, we introduce a model that generates multifractal states through geometric manipulation by placing substitutional point-like defects in a regular lattice. Specifically, we focus on the case when these defects are arranged to transform the parental lattice into a regular fractal geometry [38]. In the past few years, properties of artificial quantum fractals structures have been systematically explored both experimentally and theoretically [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. Here, our aim is to demonstrate how the self-similar structure and non-integer dimensionality of fractals offer a natural framework for designing and studying multifractal eigenstates. We consider single-particle quantum mechanics of a tight-binding model on a Sierpiński Gasket (SG), and through a comprehensive numerical analysis we demonstrate multiple crossovers between extended states and distinct stages of multifractality in the model spectrum. By calculating the fractal dimensions, we identify energy intervals where transitions to multifractal behavior occur. By examining the singularity spectrum, we gain deeper insights into the multifractal nature of individual states, unveiling a variety of behaviors including metallic and insulating NEM. In addition, we observe formation of clusters in the spatial structure of the wavefunctions that we classify as compact localized states (CLSs) [53, 54], which emerge as localized modes on the lattice subsets due to destructive interference on resonant groups of sites.
Recent advancements in spin-resolved scanning tunneling microscopy (STM) and spectroscopy techniques [55, 56, 57] have enabled precise control over quantum states in low-dimensional structures, where quantum confinement is strong. Having this in mind, we also present a detailed numerical analysis of the density-density correlation functions, which can be directly observed in such experiments. By pinpointing the regions where crossovers to NEMs occur, we compute these correlators and show that they exhibit a power-law decay combined with irregular oscillations and revivals in the frequency domain caused by self-similarity of the underlying fractal structure. Overall, these features form a robust criterion to unambiguously distinguish NEMs from both extended and localized states. Yet another verifiable aspect of multifractal systems is the mutual avoidance of eigenstates [58] at large energy separations, offering additional mechanisms by which them stand apart from quantum systems without NEMs in the spectrum.
While this study if of a purely fundamental theoretical nature, one can hope that NEMs, due to their unconventional properties, can be practically useful in certain contexts like quantum control or superconductivity. Hence, a key question here is: can multifractal states demonstrate sufficient resilience to perturbations to support their practical use? To address this, we introduce additional randomly distributed geometric defects into the system. Our analysis shows that, qualitatively, NEMs exhibit robustness in the presence of such perturbations. What is even more compelling, CLSs react to manipulation of individual atoms in a very structured and symmetric way, possibly opening an opportunity to use them for delicate control in quantum systems.
Figure 1: Transformation of regular geometry into a fractal structure: from triangular lattice (left) into a sixth generation Sierpiński gasket (right). Zoom-in: the elementary building block of the SG with a substitutional defect (yellow dot). The hopping between regular sites is $t$. The defect has on-site potential $V$, and the corresponding hoppings connecting it with the rest of the lattice are $\bar{t}\ll t\ll V.$
Results
Model
Our starting point is a uniform triangular lattice $\mathcal{L}$ containing $N$ sites, each uniquely indexed. The corresponding tight-binding Hamiltonian is a matrix of dimensions $N\times N$. Due to the uniform geometry of the lattice and assuming that the hopping occurs only between the nearest neighbors, we set the elements of the Hamiltonian to $t_{ij}=t=1$ for neighboring sites $i$ and $j$. The diagonal elements are initially set to zero, neglecting any on-site potential at this stage.
The transformation of the triangular lattice into a SG is achieved by iterative placement of defects into regions corresponding to voids in the fractal structure up to the desired generation, as shown in Fig. (1). The set of defect sites forms a subset of the original lattice denoted as $\mathcal{B}\subset\mathcal{L}$. Each defect, located at position $r_{i}$, is modeled as a delta-peaked potential: $V_{i}=V\,\delta(r-r_{i})$ with strength $V$ significantly larger than any other characteristic energy scale in the system. To make the defects act as voids of the fractal, the hoppings involving defect sites are rescaled to $\bar{t}\ll t\ll V$. The Hamiltonian reads:
[…]
We then employ exact diagonalization to determine the complete set of eigenvalues and eigenvectors of the resulting single-particle Hamiltonian, with open boundary conditions applied. For the results presented here, the interaction parameters are set to $\bar{t}=10^{-9}$, $t=1$, and $V=10^{4}$. We emphasize that this particular choice of parameters is not unique. Any set of values satisfying the inequality $\bar{t}\ll t\ll V$ is equally valid and would lead to similar qualitative outcomes.
Eigenstates statistics
The degree of delocalization in a system can be probed by examining the structure of its eigenstates in a given basis. A common approach [59, 60, 61] derived from the box-counting method [3, 18] is to analyze the asymptotic scaling properties of moments of order $q$ of the wavefunction $\Phi$ with the number of sites $N$:
$\sum_{i=1}^{N}|\langle{i}|{\Phi}\rangle|^{2q}\sim N^{-D_{q}},$ (2)
where $D_{q}$ are known as the fractal dimensions, and $\{|i\rangle\}$ form the Hilbert space basis. While Eq.(2) holds in any basis, in what follows we will be using basis of localized states, with $|i\rangle$ wavefunction being fully concentrated at site $i$. In metallic phases, eigenfunctions are extended with typical amplitude at site $i$ scaling inversely with the whole system volume. This corresponds to fractal dimension $D_{q}=1$ indicating ergodic and delocalized states. In contrast, in insulating phases, the fractal dimensions $D_{q}\rightarrow 0$, representing localized states with the amplitude being inversely proportional to the localization volume. In the critical regime, the states occupy only a vanishing fraction of the total Hilbert space [62], their fractal dimensions lie in the range $10<D_{q}<$ 1, and, moreover, exhibit non-trivial dependence on $q$, highlighting their non-ergodic, multifractal nature. In Fig. (2), we show how fractal dimension $D_{2}$ the second momentum with $q=2$ varies across the complete set of eigenstates of the tight-bidning model on the sixth generation of SG.
Figure 2: Energy spectrum of the tight-binding model on the sixth generation of the Sierpiński Gasket. The color map show the fractal dimension $D_{2}$ of each eigenstate. Multiple crossovers from extended states to NEMs are observed, with the only transition to localization occurring at $E=-2.184$.
The ground state and the first excited states have $D_{2}\geq 0.90$, which is close to the expected values for extended states in finite size systems with open boundary conditions. As energy increases, however, the majority of states display significant deviations from the typical fractal dimension values associated with extended states. Surprisingly, the system almost never enters the localized regime, with the only exception of state with energy $E=-2.184$. Pronounced deviations of $D_{2}$ from its limiting cases serve as one of indicators of the possible emergence of NEMs. They are especially large near the spectral gaps and degenerate states, where the influence of the self-similar geometry plays a key role.
The next step is to employ a more refined measure of multifractality – the singularity spectrum $f(\alpha)$, a dual measure of the fractal dimensions (see Methods) with which one can assess the strength of multifractality of a single NEM. In Fig. (3), we present some case studies. The ground state ($E=-3.666$), represented by the red curve, displays a narrow distribution peaked at $\alpha=1$, close to the shape typical for an ideal extended state, as produced, for instance, by a Gaussian Orthogonal Ensemble (thick black line). On the other hand, the singularity spectrum represented with the blue curve for $E=-2.184$ is distributed across a large domain of $\alpha$ values and has its maximum shifted far away from $\alpha=1$, as compared to the extended state. As anticipated already from its vanishingly small fractal dimension $D_{2}$, this picture closely resembles the case of an ideal localized state, with infinite width. The yellow curve depicts the singularity spectrum of the NEM degenerate state at $E=1.847$. Its parabolic shape, as well as the finite width and the apex shifted to $\alpha>1$, are in good agreement with the previous predictions [3].
Figure 3: Singularity spectrum of different states on the Sierpiński Gasket: representative cases of extended (E = -3.666, red), localized (E = -2.184, blue), and multifractal (E = 1.847, yellow) states are analyzed and compared with the Gaussian Orthogonal Ensemble (GOE), which serves as a reference for an ideal extended state. The differences in behavior can be inferred from both the position of the peak and the width of the curve.
Using these exemplary cases as a benchmark, one can examine the distribution of the singularity spectrum for different groups of eigenstates and acknowledge the emergence of different degrees of multifractality. NEMs can be not very pronounced, and their singularity spectrum can retain features of metallic or insulating phases [58, 63]. For such “mildy” multifractal states, the shift of the maximum of $f(\alpha)$ may be minor, or its parabolic shape may be distorted, along with other anomalous behaviors known as termination or freezing [3]. One distinct class of states, neither extended nor localized in a rigorous sense, is observed for certain integer energy values. Though these states have a fractal dimension typical for NEMs, their singularity spectra do not abide by the conventional classification. As one can see from Fig. (4), the left branches of the $f(\alpha)$ curves for these states are primarily concentrated around its maximum (like for the metallic states), while the right branches extend in a manner similar to that of the insulating states. This suggests that these states exhibit a form of localization, albeit not restricted to a few sites in a single region, leading us to refer to them as compact localized states (CLS) [64, 54]. The emergence of CLS can be attributed to the interplay of hidden symmetries and destructive interference at resonant sites located along the reflection axes—characteristics inherent to the underlying fractal geometry, as visualized in the amplitude profiles (see the Supplemental Material). In fractal geometries, CLSs were first identified and numerically characterized in [65].
Other examples or multifractal states are presented in the Supplemental Material. Some of them have spatial structures analogous to the bulk and edge modes observed in fractal geometries [45, 50].
Figure 4: Singularity spectrum for the Sierpiński Gasket: representative cases of compact localized states. In each case, the left side of the $f(\alpha)$ curve clusters around its peak, while the right side extends in a way similar to insulating states. This unique behavior blends features of both extended and localized states, indicating a novel hybrid phase.
We further expand upon our findings by probing the density-density correlators that encode for the overlap between eigenstates and serve as an additional tool to discriminate the NEMs, particularly at critical points when the correlator $\mathcal{F}(\omega,E)$ decays as a power law (see Methods). Fig. (5) illustrates the behavior of $\mathcal{F}(\omega,E)$ for the previously selected representatives of ergodic, localized and NEM states. As predicted by theory, the two-points correlator for the extended (red curve) and localized (blue curve) states exhibits a broad plateau across several orders of $\omega$. The intensity of the correlator reflects the degree of eigenstates overlap: it is low for extended states that share minimal information across a wide spatial region, and high for localized states that are concentrated in a confined region. However, in this particular case of localization, the plateau is especially pronounced, and from the amplitude profile of the state (see the Supplemental Material), it is evident that localization occurs within a single region, though with a slower decay. This suggests that localization is algebraic, rather than of the Anderson type. At large values of energy separation $\omega$, the correlator for both extended and localized states approaches the uncorrelated regime, $\mathcal{F}(\omega,E)=1$, as marked by the typical GOE statistics. The NEM state (yellow curve) qualitatively behaves as expected, displaying a shorter plateau followed by a power-law decay towards the uncorrelated regime as $\mathcal{F}(\omega,E)\sim\left(\omega/E_{Th}\right)^{-\mu}$. The exponent $\mu$ and the Thouless energy $E_{Th}$ are model-dependent and, unlike in disordered quantum systems, the lack of variable parameters in our Hamiltonian makes a possible analytic fit difficult. Notably, the density-density correlator behavior at higher frequencies is marked by pronounced oscillations with clear revivals and dips below the uncorrelated limit. This non-trivial behavior signals the presence of anomalies, reminiscent of effects observed in certain disordered quantum systems stemming from random matrix theory [66, 67, 68]. In our model, these oscillations arise from the competition between the eigenstate avoidance and the resonant states interference driven by the self-similar structure of the Sierpinski Gasket (SG). As shown in [58], the eigenstates avoidance is responsible for dips in $\mathcal{F}(\omega,E)$ falling below the uncorrelated limit. Such dips indicate that, when a state predominantly occupies a site $r$, this site is likely to be vacant for other states: a behavior that is known as eigenstates stratification. On the other hand, revivals are not typical for randomly disordered systems and specific only for the considered case of regular fractal geometry.
Given that genuine multifractality should persist across different scales, we examine the scaling of correlator $\mathcal{F}(\omega,E)$ with respect to system size $N$ (see Fig. (6)) by computing it for subsequent generations of SG. The rescaled correlator $\mathcal{F}(\omega,E)/N^{(1-D_{2})}$ for the multifractal state at $E=1.847$, shows remarkable consistency across different generations of the SG. The extent of the plateau and the locations of peaks and dips remain unchanged. Aside from minor finite-size corrections affecting the curve heights, the overall qualitative agreement is nearly perfect.
Figure 5: Frequency correlator $\mathcal{F}(\omega,E)$ for three characteristic cases: extended (E = -3.666, red), localized (E = -2.184, blue), and multifractal (E = 1.847, yellow). The ideal case of uncorrelated systems as described by the Gaussian Orthogonal Ensemble (GOE) is represented by the dashed black line. A broad plateau at small values of energy separation $\omega$ is a hallmark of both extended and localized states, indicating non-vanishing correlations. In contrast, NEMs exhibit only a narrow plateau, which diminishes to a singular point in the thermodynamic limit, followed by a distinct power-law decay. At larger values of $\omega$, only NEMs display revivals of the correlator, alongside a stratification phenomenon (in the Supplemental Material, examples with more pronounced revivals are presented). According to theory, the transition between phases at small and large energy separations should occur smoothly. However, in our simulations, discontinuities can appear in the correlator for both extended and localized states within the interval $1010^{-1}<\omega<$ 10 due to numerically zero denominators in the computation of the density of states. Lines are guide to eye.
Figure 6: Rescaled frequency correlator $\mathcal{F}(\omega,E)/N^{(1-D_{2})}$ for the NEM at energy $E=1.847$ shown for different generations of the Sierpiński Gasket. The Gaussian Orthogonal Ensemble, dashed black line, serves as the reference to an ideal uncorrelated system. The curves corresponding to different generations exhibit consistent behavior highlighting the scale-invariant properties of the NEM state. Lines are guide to eye.
Robustness and quantum control
So far, we have considered the case of a clear fractal structure. In reality, random defects existing on top of the regular fractal lattice will be likely unavoidable. Damaging the regular self-similar pattern of the fractal structure could trivially drive the NEMs of the original system towards a localized phase, and this needs to be analyzed.
In Fig. (7), we show the fractal dimension $D_{2}$ for the set of eigenstates of the SG lattice in the presence of a number of additional random defects. One can see a clear general tendency for the NEMs to retain their multifractal nature, with $D_{2}$ remaining distinct from the limiting case of localized states even for a large number of introduced defects. While individual NEMs may not exhibit robustness to disorder in a strict sense, the latent symmetries of the SG hierarchical structure act to prevent a full transition to localization, preserving the overall multifractal regime. The closing of gaps and the lifting of degeneracies reduce the average fractal dimension, yet the eigenstates of the system continuously reorganize themselves through resonances and stratification. This reorganization enables the NEMs to adapt, exploiting new patterns of multifractality that persist across different scales. Upon the introduction of additional defects, a broader spectrum of multifractal behaviors is observed, ranging from extended to insulating multifractal phases before an eventual transition to localization becomes inevitable.
Figure 7: Energy spectrum of the single-particle tight-binding model on the sixth generation of SG with additional random defects (ARD). Configurations with 50, 100, 150 and 200 ARD are analysed. The color indicates the fractal dimension $D_{2}$ of each eigenstate. In the lower right corner of each subplot, the average of the fractal dimensions $\langle D_{2}\rangle$ over the spectrum is displayed, quantifying the tendency to localization as the number of defects increase.
The overall tendency of NEMs to persist upon introducing random defects can be complemented with a more detailed analysis of spatial structure of a specific state. In particular, compact localized states (CLSs) are of special interest due to their promise for applications to quantum information storage and transfer [69, 64]. As an illustrative example, we consider a state hosted by the sixth generation the SG lattice, Fig. (8), that can be created using atomic quantum simulators [70, 57]. By tuning the voltage to align the energy level with the CLS formation at $E=1.000$, one can observe a nine-fold degenerate CLS exhibiting non-zero wavefunction amplitude solely in several isolated regions adjacent to the internal edges of the fourth level of the SG, Fig. (8a). The isolated clusters of the wavefunction intensity of CLSs remain unchanged unless perturbations directly affect the specific lattice sites on which these clusters have their support. At the same time, if defects are systematically introduced on sites within each cluster, the degeneracy of the state is lifted, and the wavefunction amplitude is redistributed among the remaining clusters without involving the other non-resonant sites Fig. (8b-d), allowing for targeted controlled manipulation of the wavefunction profile and offering a potential for accurate quantum control. As we show in the Supplemental Material, even if a larger number of random defects are introduced, and all the major clusters are destroyed, these disruptions lead to formation of new quantum interference patterns, wherein the information from a destroyed cluster stabilizes around lower levels of hierarchy of the fractal, which again results in formation of CLSs.
Figure 8: Amplitude profile of the compact localised state at energy level $E=1.000$. Profile (a) in the original model without defects, and (b-d) with additional defects (marked by pink crosses) placed at three, six, and eight targeted sites, respectively. Each defect, positioned within its respective cluster, completely blocks information propagation in that region, while causing lossless redistribution of information among the remaining clusters. The colorbar indicates of $|psi|$, with red corresponding to high amplitude values and black to nearly zero ones.
Discussion
With this work, we continue the research line of studying quantum phenomena in artificial fractal geometries, specifically focusing on the case of the Sierpiński gasket. We have shown that in clean fractal, without any added randomness, non-ergodic multifractal states naturally emerge, complementing the prior studies on wavefunction localization [45], quantum transport [21], and eigenstate statistics in fractals, and fractal energy spectra [71]. We performed a multi-tool analysis of both the overall properties of SG spectrum and the specific features of individual NEMs and provided sensitive indicators to distinguish between different types of them. Among other findings, we demonstrated that one particular type of states – compact localized states – respond in a highly symmetric and structured way to introducing random disorder, making them an interesting platform for quantum control.
Not to limit ourselves with only one fractal geometry, in the Supplemental Material, we generalized our findings to another structure: the negative Vicsek fractal, where vacancies (rather than the remaining lattice sites) form a fractal of rather low Hausdorff dimension ($\simeq 1.46$), and next-nearest hoppings are allowed. The main goal to consider this particular example was that it creation requires introducing a minimal number of defects as compared with structures like SG or the Sierpiński carpet, making its experimental realization easier. While most of the states in this geometry are ergodic, NEMs still emerge in certain parts of the spectrum. Besides, it demonstrates the emergence of non-ergodic multifractal states (NEMs) in a system with beyond nearest-neighbor kinetics [15] and large translationally invariant clean sub-domains.
We suggest that artificial fractal structures – systems that have already been created in a lab – are a natural platform to create and study multifractal states experimentally, which so far – to our knowledge – has been possible only in doped semiconductors [72] and yet remains an ongoing challenge. Our numerical results, particularly on the two-point density-density correlation function $\mathcal{F}(\omega,E)$, provide a direct method to distinguish NEMs using scanning probe microscopy techniques (recent research has also demonstrated applicability of similar methods in acoustic metamaterials [73] and photonic systems [51, 74, 75]).
The signatures of multifractal states that we have analyzed here, lay a natural ground for further research. First of all, we have considered only single-particle quantum mechanics. Further studies could focus on how the observed phenomena change upon introducing interactions, potentially relating to such phenomena as many-body localization, Hilbert space fragmentation [27], and quantum scars [76, 77]. The demonstrated properties of compact localized states, also known for their ability to support dispersionless wave propagation [78, 64], can be deeper analyzed in connection with controlled information flow between targeted regions of the system, offering new approaches to quantum control.
Methods
Singularity spectrum
The generalized fractal dimension $D_{q}$ of eigenstates $\psi_{i}$ with energy E is defined as [18]:
$D_{q}(E)=\frac{1}{\left(1-q\right)\ln{N}}\ln{\sum_{i=1}^{N}\sum_{r=1}^{N}|\psi_{i}(r)|^{2q}\,\delta\!\left(E-\lambda_{i}\right)}$ (3)
The information encoded in the fractal dimensions can be refined by analysing the singularity spectrum $f(\alpha)$ [79, 25, 80], which, for each $\alpha$, can be regarded as a fractal dimension of the set of points, where the wavefunction intensity is $|\psi|^{2}\sim N^{-\alpha}$ [3]. It can be obtained as the Legendre transform of the mass exponent dimensions $\tau_{q}\equiv(q-1)\,D_{q}$:
$\alpha_{q}$ $=\frac{d\tau_{q}}{dq},\quad\quad q=\frac{df}{d\alpha}$
$f(\alpha_{q})=\alpha_{q}q-\tau_{q}.$ (4)
The maximal value of this function is always $f(\alpha_{max})=1$, and it obeys the relation: $f(1+\alpha)=f(1-\alpha)+\alpha$ [81]. For ergodic states, the singularity spectrum shows a narrow distribution around the maximum, converging to a $\delta(\alpha-1)$ in the limit of ideal extended states, as it happens, for example, in the case of Gaussian Orthogonal Ensemble. For ideally localized states, $f(\alpha)$ has its support in just two points, $\alpha=0$ with $f(0)=0$, and $\alpha=\infty$ with $f(\infty)=1$. For multifractal states, a smooth parabolic spectrum is typically expected, which should be independent of the size of the system. However, the degree of multifractality can significantly influence the result, affecting both convexity of the spectrum and its support range. Properties of the singularity spectrum can be analyzed in more detail through extrapolation to the thermodynamic limit [36, 82] or using graphical methods [68], but we do not employ these approaches here.
Two-eigenstates correlator
Following Berezinskii and Gorkov [83], we define the two-point density correlator at energy $E$ for an arbitrary pair of points $(r,r^{\prime})$ as:
$\rho_{E}(r)\rho_{E+\omega}(r^{\prime})=\dfrac{1}{n(E)}\sum_{i,j}|\psi_{i}(r)|^{2}|\psi_{j}(r^{\prime})|^{2}\delta(E-\lambda_{i})\delta(E-\lambda_{j}+\omega),$ (5)
where $n(E)$ is the density of states, and $\omega$ is the energy difference between two states. In the case of extended states, the corresponding wavefunctions are spread over all $N$ sites of the system (here, we always deal with fractals embedded in $2d$ space), and each term in the sum scales as $N^{-2}$, vanishing in the limit $N\rightarrow\infty$. In the localized case, the wavefunctions have their support in a finite volume proportional to square of the characteristic localization length, and the density correlator is dominated – apart from a nonsingular contribution – by a pole singularity:
$\rho_{E}(r)\rho_{E+\omega}(r^{\prime})\sim A_{E}(r-r^{\prime})\delta(\omega),$ (6)
where the spectral function
$A_{E}(r-r^{\prime})=\dfrac{1}{n(E)}\sum_{i}|\psi_{i}(r)|^{2}|\psi_{i}(r^{\prime})|^{2}\delta(E-\lambda_{i})$ (7)
is directly related to the number of sites in the support of the considered quantum states and can be linked to the fractal dimension $D_{q}$, see Eq.(3).
This probe has been widely used to study translationally invariant disordered quantum systems [58, 80]. For the fractal geometries that lack translational invariance, it is more handy to work with the spectral correlation function $\mathcal{F}(\omega,E)$ defined as:
$\mathcal{F}(\omega,E)=\frac{N\sum_{r}\sum_{i,j}|\psi_{i}(r)|^{2}|\psi_{j}(r)|^{2}\,\delta\!\left(E-\lambda_{i}-\dfrac{\omega}{2}\right)\delta\!\left(E-\lambda_{j}+\dfrac{\omega}{2}\right)}{\sum_{i,j}\delta\!\left(E-\lambda_{i}-\dfrac{\omega}{2}\right)\delta\!\left(E-\lambda_{j}+\dfrac{\omega}{2}\right)}.$ (8)
This function effectively quantifies overlaps between eigenstates separated by energy $\omega$. To perform numerical simulations, we smeared the $\delta$-functions with Gaussian distributions.
For ideal extended states, such as those described by the Gaussian Orthogonal Ensemble (GOE), the correlator $\mathcal{F}(\omega,E)=1$ across the entire spectral range, and the eigenfunctions are completely uncorrelated [58, 84]. Analysing more realistic scenarios, such as typical metals, a broad plateau appears at $\mathcal{F}(\omega,E)>1$, and its height increases as the states become more localized. As the separation between eigenstates grows, the plateau region becomes bounded by a cutoff energy – the Thouless energy $E_{Th}$ – at the scale of the lattice spacing [85, 86], beyond which the system returns to the uncorrelated regime. For a multifractal state, the plateau is present only within a narrow energy interval, and the Thouless energy shrinks to zero in the thermodynamic limit. In this case, the uncorrelated limit at large values of $\omega$ is preceded by power-law decay $\mathcal{F}(\omega,E)\sim\left(\omega/E_{Th}\right)^{-\mu}$ in the intermediate range of energies, where $\mu$ is the critical exponent [85]. With system size $N$, wavefunction amplitudes on each site scale as $N^{-D_{2}}$. Consequently, the height of the plateau at low energies scales as $\mathcal{F}(0,E)\propto N^{1-D_{2}}$. For NEMs, this implies that $\mathcal{F}(\omega,E)$ for different system sizes $N$ should collapse onto a single curve when the vertical axis is rescaled by $N^{1-D_{2}}$.
Availability of data
Data are available upon reasonable request.
Acknowledgements
We thank C. Morais Smith, M.R. Slot, A. Moustaj, M.Conte and Y. in ’t Veld for the insightful discussions.
This work was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program, grant agreement 854843-FASTCORR. The data that support the findings of this study are available from the corresponding author upon reasonable request.